# A Module-theoretic Introduction to Abstract Elementary Classes

### v1

@@ 2.4. First results
Fact 2.9 (Shelah’s Presentation Theorem, [She87, Lemma 1.8]).
@@ 2.5. Basic Properties
The axioms of Abstract Elementary Classes are very bare bones. One the one hand, this is desirable because itt makes it much easier to determine which classes are AECs. On the other hand, it does not give much to use to prove theorems. There is a common set of structural properties that have emerged as common hypotheses in theorems, and lemmas to prove when examining specific classes. As mentioned, we listed the relevant properties earlier when going through examples to streamline the presentation.
@@
    $\mathbb{K}$ has the joint embedding property iff for every $M_{0},M_{1}\in\mathbb{K}$, there is $N\in\mathbb{K}$ and $f_{\ell}:M_{0}\to N$ for $\ell=0,1$. In pictures,
@@
    Interestingly (to those looking for more set theory), the converse fails very bad without joint embedding. It is well-known that there is a computable Hanf number for arbitrarily large models, where an AEC $\mathbb{K}$ has arbitrarily large models iff it has models of size at least $\beth_{(2^{\text{LS}(\mathbb{K})})^{+}}$. On the other hand, one can construct an AEC with arbitrarily large models that has no maximal models above a measurable cardinal (if it exists), and Baldwin-Shelah [BS] have constructed such an example from a complete $\mathbb{L}_{\omega_{1},\omega}$ sentence.
@@
We can use these properties to build a monster model (Definition 2.13). Before we do this, we introduce an important notion in Abstract Elementary Classes: Galois types. The strength of monster models is best understood by how they simplify this notion. For those familiar with first-order model theory, a Galois type is a semantic version of the normal syntactic type101010‘Type’ is best understood here as a synonym of ‘description.’ that describes the class of elements an element could be mapped to.
@@
    where $N^{\prime}$ is some/any model $M_{0}\prec_{\mathbb{K}}N^{\prime}\prec_{\mathbb{K}}N$ so $a\in N^{\prime}$ and $\|N^{\prime}\|=\|N\|$.
@@
We have implicitly used the basic proposition that atomic equivalence is reflexive and symmetric, so the transitive closure of $\equiv_{at}$ is an equivalence relation. More importantly, if $\mathbb{K}$ satisfies amalgamation, then $\equiv_{at}$ is already transitive (see [Bal09, Exercise 8.8]). The existence of a monster model is often-hand waved using ‘inessential’ large cardinals or other metamathematical tricks (see [HK23]), but we follow [Gro1X] to make this precise inside of ZFC.
@@
    $\text{gtp}(a/M;\mathfrak{C})=\text{gtp}(b/M;\mathfrak{C})\iff\exists f\in\text{Aut}_{M}\mathfrak{C}\text{ so }f(a)=b$
@@
Note that $\mathfrak{C}$ will be much larger than $\theta$ unless $\theta$ is an inaccessible cardinal.
@@
One of the key properties of syntactic types in first-order logic follows from the facts that they are sets of formulas and each formula has finitely many parameters. This immediately gives a local character (even finite character) to syntactic types: any two different syntactic types differ on some finite subset of their domain. With Galois types having a semantic definition (as orbits in monster models or equivalence classes of amalgams), there is no reason to expect any local character to hold. However, Grossberg and VanDieren [GV06b, GV06c, GV06a] isolated a local character property of Galois types–called tameness–and used it to prove a version of Shelah’s Categoricity Conjecture (see Section 3.1.1) under the additional hypothesis of a monster model. Since then, tameness has emerged as a very powerful assumption to make and a property that holds in many natural examples; see the survey [BV17b] for more (and see the instances of this property listed in Example 2.7). As above, there are many parameterizations and variants of this property that we suppress (but are listed in the survey).
@@
Although Galois types are defined semantically, in nice cases one can often prove that there are syntactic (see Fact 3.14 for a partial list). In this case, the class is $<\omega$-tame for free!
We finish we two properties that are common in algebraic examples: universal classes and closure under intersections.
@@
Recall from the introduction that $(\mathbb{K}^{\text{R-Mod}},\subset)$ is a universal Abstract Elementary Class and $(\mathbb{K}^{\text{R-Mod}},\subset_{\text{pr}})$ is an Abstract Elementary Class with intersections. Both of these properties greatly simplify the conditions necessary to check equality of Galois types wether or not amalgamation holds.
Fact 2.17.
Let $\mathbb{K}$ be an Abstract Elementary class, $M\in\mathbb{K}$, and $\text{gtp}(a/M;N),\text{gtp}(b/M;N^{\prime})\in\text{S}(M)$.
@@
    $\text{gtp}(a/M;N)=\text{gtp}(b/M;N^{\prime})\iff\exists f:\langle aM\rangle_{N}\cong_{M}\langle aM\rangle_{N}\text{ s.t. }f(a)=b$
@@ 3.1.1. Categoricity transfer in abstract Abstract Elementary Classes
The behavior of ‘any two structures of the same size being isomorphic’ is called categoricity in power131313‘Power’ is an old-fashioned term for size or cardinality. Also, the term ‘categoricity’ once meant that every structure in a class was isomorphic, in the sense that all of the models were required to be isomorphic regardless of size. However, in the context of model theory, this is very uncommon, so ‘categoricity’ with no modification is normally used to mean ‘categoricity in power.’.
In 1954, Łoś observed that, among countable first-order theories, there tended to be an all or nothing behavior with categoricity in power, and conjectured for the class of models of a countable first order theory, either the theory was categorical in every uncountable cardinal or it was categorical in no categorical cardinals; the question of countable categoricity turns out to be independent from the rest.
Morley [Mor65] proved this in the affirmative in his 1964 PhD thesis. This was followed by work of Baldwin and Lachlan [BL71] that showed that this result can be explained by finding—in a class of models of a countable theory that is categorical in an uncountable cardinal—admit a dimension invariant that determines the isomorphism class of the model. Shelah [She74] further extended Morley’s result showing that it also holds amongst uncountable theories, although the cutoff must be larger.
@@
Fact 3.1 (Morley [Mor65], Baldwin-Lachlan [BL71], Shelah [She74]).
@@
In particular, $\mu_{\lambda}=\lambda^{+}$. If $\lambda=\omega$, then there is a dimension associated with the models of $T$ that determines the isomorphism class (including for countable models). [^14] Unfortunately, Shelah’s extension to uncountable theories does not come with an associated dimension. This means that the motivating example from real vector spaces is not actually explainable by the abstract model theory. However, the linear dimension for vector spaces of countable field (such as $\mathbb{Q}$) or the transcendence dimension for algebraically closed fields of a fixed characteristic is discoverable via the Baldwin-Lachlan result).
@@
Conjecture 3.2 (Shelah’s Categoricity Conjecture).
@@
The expectation is that $\mu_{\lambda}=\beth_{\left(2^{\lambda}\right)^{+}}$, but even the more general eventual version stated above would be a huge advantadge.
@@
    Suppose that $2^{\kappa^{+n}}<2^{\kappa^{+(n+1)}}$ for all $n<\omega$. If $\mathbb{K}$ is categorical in $\kappa^{n}$ for all $n<\omega$, then $\mathbb{K}$ is categorical in all $\kappa^{\prime}\geq\kappa$ ([She09c, Conclusion 12.43])
@@
    Suppose $\mathbb{K}$ has amalgamation, no maximal models and is $\chi$-tame. Set $\mu_{\lambda}=\lambda^{+}+\kappa$. If $\mathbb{K}$ is categorical in $\kappa^{+}>\mu_{\lambda}$, then $\mathbb{K}$ is categorical in all $\kappa^{\prime}>\kappa^{+}$. ([GV06a, Theorem 5.2], Upward Categoricity from a Successor with Tameness)
@@
We should also comment on the description of Shelah’s Categoricity Conjecture as a ‘test question.’. This is used to indicate that it is not expected that a positive solution will lead to new results by citing the true statement. Instead, proving the conjecture is a test of the robustness of the structure theory of Abstract Elementary Classes, and it is the techniques used to solve it that will be used to prove other theorems. This has been scene in the partial results listed above, e.g., tameness in Fact 3.3.(3) or good $\lambda$-frames in Fact 3.3.(2).
@@ 3.1.2. Categoricity in Abstract Elementary Classes of Modules
We now situate categoricity in the study of Abstract Elementary Classes of modules. An important point is that categoricity is solely a property of the models in an Abstract Elementary Class $\mathbb{K}$ and is not impacted by the particular choice of strong substrcuture $\prec_{\mathbb{K}}$ (although some strong substructure must be given to witness that the collection of modules form an Abstract Elementary Classes). One theme throughout the results is tha the threshold cardinals (which we normally denote $\mu_{\lambda}$) are much lower in the module examples than in general.
@@
Given a module $M$ and a cardinal $\lambda$, we use $M^{(\lambda)}$ to denote the $\lambda$-fold direct sum of $M$, $\oplus_{\lambda}M$.
@@
In nice classes, there is a tight relation between categoricity and the number of strongly indecomposable modules.
@@
There are many examples of classes of modules that turn out to be nice. In many cases, the categorical side of dichotomoy (where $\mathbb{K}$ is categorical in some/all $\lambda>\mu_{\mathbb{K}}$) also comes with a characterization of the ring $R$. The various definitions of the ring properties are not crucial for our purposes, but are available at the citations.
@@
    $\mathbb{K}^{\text{R-Mod}}$ is categorical in some power iff $R$ is the ring of matrices over some division ring. [MA23a, Lemma 2.12]
@@
    This class is categorical in some power iff $R$ is left pure semi-simple and, therefore, all modules are locally pure-injective (and this reduces to the previous case). [MA23a, Theorem 3.4]
@@
    This class is categorical in some power iff $R$ is left Noetherian. [MA23a, Theorem 3.11]
@@
    This class is categorical in some power iff $R$ is left Noetherian. [MA23a, Theorem 3.19]
@@
    This class is categorical in some power iff $R$ is left perfect and $R/J(R)$ is the ring of matrices over some division ring iff $R$ is the ring of matrices over a local ring with left $T$-nilpotent maximal ideal.
The extra information about $R$ in these prongs comes from showing that relevant properties interact nicely with direct sums and any special module we care about is the direct sum of our strongly indecomposable generator. We give an example below. Recall that if every absolutely pure left $R$-module is injective, then $R$ is left Noetherian.
@@
Another direction [vT, Trl] for categoricity focuses on strong splitters as the special models in the context of deconstructible classes of modules.
@@
    The class $\mathbb{K}$ is deconstructible iff there is some $\kappa$ such that every $M\in\mathbb{K}$ has a $\mathbb{K}_{*<\kappa}$-filtration, where $\mathbb{K}_{*<\kappa}$ are the $<\kappa$-presented modules in $\mathbb{K}$.
@@
    for some $\kappa$ so $M$ is $\leq\kappa$-presented. [^16] Note the definition requires this condition for all direct sums, but [Trl, Lemma 2.2] shows that this single set is enough..
Note that $M$ being a strong splitter means that given any $N\supset M$ such that $N/M\cong\oplus_{\kappa}M$, we must have $N\cong\oplus_{\kappa+1}M$.
Note that not all deconstructible classes of modules are Abstract Elementary Classes (implicitly using the submodule relation unless we state otherwise) and the following results cover all deconstructible classes. However, there are many examples of deconstructible classes that are Abstract Elementary Classes: any Roots of Ext (Example 2.7.(7)) that form an AEC are deconstructible [vT, Theorem 2.2].
In deconstructible classes, strong splitters are closely tied to categoricity. In fact, the existence of a single non-strong splitter massively restricts the categoricity spectrum.. A helpful observation is that it suffices to check a particular model to see if there are any non-strong splitters.
@@
Fix a deconstructible class $\mathbb{K}$ generated by $\mathcal{C}$. Define
@@
Let $\mathbb{K}$ be a deconstructible class of $R$-modules. Suppose there is a non-strong splitter $M\in\mathbb{K}$ of size $\kappa$. Then $\mathbb{K}$ is categorical in at most one cardinal $\lambda\geq\kappa$ and such a $\lambda$ is either equal to $\kappa$ or singular.
Of course, these restrictions on $\lambda$ beg the question if such a $\lambda$ could possibly exist. This question is open [Trl, p. 381].
@@
To do this, build a continuous, increasing chain $\{M_{\alpha}:\alpha<\mu\}$ starting with $M_{0}=M^{(\lambda)}$ and using the following successor step: if $M_{\alpha}$ is isomorphic to some $\oplus_{\chi}M$ (necessarily with $\chi\geq\lambda$), then the definition of not strongly splitting means that there is an extension $M_{\alpha+1}\supset M_{\alpha}$ with $M_{\alpha+1}/M_{\alpha}\cong M$ and $M_{\alpha+1}\not\cong M^{(\chi+1)}$. The first condition ensures $M_{\alpha+1}$ (and the eventual union) is in the deconstructible class, and the second condition will force us away from the union being a direct sum itself.
Set $M^{*}_{\mu}=\bigcup_{\alpha<\mu}M_{\alpha}$. The proof that $M^{*}_{\mu}$ is not isomorphic to $\oplus_{\mu}M$ is different in each case, but the essential argument is to a back and forth argument from a supposed bijection that contradicts our construction: it finds a stage $M_{\beta}\subset M_{\beta+1}$ such that $M_{\beta}$ is a direct sum of $M$ and $M_{\beta+1}=M_{\beta}\times M$. †
On the structure side, we may assume that all models are strong splitters. Then categoricity reduces to finding a power will all direct sums of generating modules to that power are isomorphic.
@@
Second, assume the right side and some $N_{0}\in\mathcal{C}$. Fix $M\in\mathbb{K}_{\lambda}$ and we will show this is isomorphic to $N_{0}^{(\lambda)}$. Since every module in $\mathcal{C}$ is a strong splitter, we have that
@@
$M$ $=$ $M_{0}\oplus M_{1}$
For $M_{0}$, we have $|\mathcal{C}_{0}|\leq|\mathcal{C}|<\lambda$, so $\|M_{0}\|<\lambda$. For $M_{1}$, we use our assumption and the fact $\kappa_{N}\geq\nu$ to conclude
@@
By our assumption, $M_{1}$ is isomorphic to a direct summand of $N_{0}^{(\lambda)}$ and $M\cong M_{1}\oplus N_{0}^{(\lambda)}$. So Eilenberg’s trick can be adapted (see the proof of [vT, Lemma 4.5]) to show

### v2

@@ 2.4. First results
Fact 2.10 (Shelah’s Presentation Theorem, [She87, Lemma 1.8]).
@@
Another basic result in Abstract Elementary Classes is the computation of Hanf numbers (for model existence). The question at hand is whether an Abstract Elementary Class must have arbitrarily large models. In first-order model theory, the Compactness Theorem (2.7) implies that any elementary class axiomatized by $T$ with a model of size at least $|T|+\aleph_{0}$ has models of arbitrarily large sizes (this means the Hanf number of such classses is $|T|+\aleph_{0}$). Without that theorem, the situation in Abstract Elementary Classes is much more delicate.
Fixing a $\lambda$ to be the $\text{LS}(\mathbb{K})$ cardinal, an application of the Axiom of Replacement implies there is a cardinal $H(\lambda)$ such, for any Abstract Elementary Class $\mathbb{K}$ with $\text{LS}(\mathbb{K})\leq\lambda$, if $\mathbb{K}$ has a model of size at least $H(\lambda)$, then it has arbitrarily large models (Hanf [Han] was the first to make this kind of argument, but not in the context of Abstract Elementary Classes). This argument shows that such a cardinal exists, but is no help in computing it. However, for Abstract Elementary Classes, a mostly optimal bound is known:
Fact 2.11.
Given a cardinal $\lambda$ and an Abstract Elementary Class $\mathbb{K}$ with $\text{LS}(\mathbb{K})\leq\lambda$, if $\mathbb{K}$ has a model of size $\geq\beth_{(2^{\lambda})^{+}}$ (or even just models of sizes cofinal in this cardinal), then $\mathbb{K}$ has models of all sizes $\geq\text{LS}(\mathbb{K})$.
Thus, $H(\lambda)\leq\beth_{(2^{\lambda})^{+}}$.
The proof of this first uses Shelah’s Presentation Theorem to show that ‘the Hanf number of Abstract Elementary Classes with $LS(\mathbb{K})\leq\lambda$’ is the same as ‘the Hanf number of classes axiomatized by theories in $\mathbb{L}_{\lambda^{+},\omega}$.’. Then, this second value is known classically by using Morley’s Omitting Types Theorem (due to Morley [Mor65b] and Chang [Cha68], see [Bal09, Theorem A.3] for a proof).
@@ 2.5. Basic Properties
The axioms of Abstract Elementary Classes are very bare bones. On the one hand, this is desirable because this simplicity makes it much easier to determine which classes are AECs. On the other hand, the minimalist assumptions make it hard (or impossible) to prove deep theorems. There is a common set of structural properties that have emerged as common hypotheses in theorems, and lemmas to prove when examining specific classes. As mentioned, we listed the relevant properties earlier when going through examples to streamline the presentation.
@@
    $\mathbb{K}$ has the disjoint amalgamation property iff for every span as above, we can find an upper bound $N$ as there with the additional property that
    $f_{1}[M_{1}]\cap f_{2}[M_{2}]=f_{1}[M_{0}]$
    That is, the image of $M_{1}$ and $M_{2}$ are disjoint except for the required overlap on the image of $M_{0}$.
  3. (3)
    $\mathbb{K}$ has the joint embedding property iff for every $M_{0},M_{1}\in\mathbb{K}$, there is $N\in\mathbb{K}$ and $f_{\ell}:M_{\ell}\to N$ for $\ell=0,1$. In pictures,
@@
    Interestingly (to those looking for more set theory), the converse fails very badly without joint embedding. Recalling Fact 2.11, we know an AEC $\mathbb{K}$ has arbitrarily large models iff it has models of size at least $\beth_{(2^{\text{LS}(\mathbb{K})})^{+}}$. On the other hand, one can construct an AEC with arbitrarily large models that has no maximal models above a measurable cardinal (if it exists), and Baldwin-Shelah [BS] have constructed such an example axiomatized by a complete $\mathbb{L}_{\omega_{1},\omega}$ sentence.
@@
We can use these properties to build a monster model (Definition 2.15). Before we do this, we introduce an important notion in Abstract Elementary Classes: Galois types. The strength of monster models is best understood by how they simplify this notion. For those familiar with first-order model theory, a Galois type is a semantic version of the normal syntactic type111111‘Type’ is best understood here as a synonym of ‘description.’ that describes the class of elements an element could be mapped to.
@@
[…]
@@
We have implicitly used the basic proposition that atomic equivalence is reflexive and symmetric, so the transitive closure of $\equiv_{at}$ is an equivalence relation. More importantly, if $\mathbb{K}$ satisfies amalgamation, then $\equiv_{at}$ is already transitive (see [Bal09, Exercise 8.8]). The existence of a monster model is often hand-waved using ‘inessential’ large cardinals or other metamathematical tricks (see [HK23]), but we follow [Gro1X] or [Bon, Theorem 3.21] to make this precise inside of ZFC.
@@
    $\text{gtp}(a/M;\mathfrak{C})=\text{gtp}(b/M;\mathfrak{C})\iff\exists f\in\text{Aut}_{M}\mathfrak{C}\text{ s.t. }f(a)=b$
@@
Note that a $\theta$-monster model $\mathfrak{C}$ will be much larger than $\theta$ unless $\theta$ is an inaccessible cardinal.
@@
One of the key properties of syntactic types in first-order logic follows from the facts that they are sets of formulas and each formula has finitely many parameters. This immediately gives a local character (even finite character) to syntactic types: any two different syntactic types differ on some finite subset of their domain. With Galois types having a semantic definition (as orbits in monster models or equivalence classes of amalgams), there is no reason to expect any local character to hold. However, Grossberg and VanDieren [GV06b, GV06c, GV06a] isolated a local character property of Galois types–called tameness–and used it to prove a version of Shelah’s Categoricity Conjecture (see Section 3.1.1) under the additional hypothesis of a monster model. Since then, tameness has emerged as a very powerful assumption to make and a property that holds in many algebraic examples; see the survey [BV17b] for more (and see the ubiquity of tameness in the classes of Example 2.8). As above, there are many parameterizations and variants of this property that we suppress (but are listed in the survey).
@@
Although Galois types are defined semantically, in nice cases, one can often prove that they are syntactic (see Fact 3.14 for a partial list). In this case, the class is $<\omega$-tame for free!
We finish with two properties that are common in algebraic examples: universal classes and closure under intersections.
@@
Recall from the introduction that $(\mathbb{K}^{\text{R-Mod}},\subset)$ is a universal Abstract Elementary Class. These two properties greatly simplify the conditions necessary to check equality of Galois types whether or not amalgamation holds.
Fact 2.19.
Let $\mathbb{K}$ be an Abstract Elementary class, $M\in\mathbb{K}$, and $\text{gtp}(a/M;N),\text{gtp}(b/M;N^{\prime})\in\text{S}_{\mathbb{K}}(M)$.
@@
    $\text{gtp}(a/M;N)=\text{gtp}(b/M;N^{\prime})\iff\exists f:\langle aM\rangle_{N}\cong_{M}\langle bM\rangle_{N}\text{ s.t. }f(a)=b$
@@ 3.1.1. Categoricity transfer in abstract Abstract Elementary Classes
The behavior of ‘any two structures of the same size being isomorphic’ is called categoricity in power141414‘Power’ is an old-fashioned term for size or cardinality. Also, the term ‘categoricity’ once meant that every structure in a class was isomorphic, in the sense that all of the models were required to be isomorphic regardless of size. However, in the context of model theory, this is very uncommon, so ‘categoricity’ with no modification is normally used to mean ‘categoricity in power.’.
In 1954, Łoś observed that, among countable first-order theories, there tended to be an all or nothing behavior with categoricity in power, and conjectured for the class of models of a countable first order theory, either the theory was categorical in every uncountable cardinal or it was categorical in no uncountable cardinals; the question of countable categoricity turns out to be independent from the rest.
Morley [Mor65a] proved this in the affirmative in his 1964 PhD thesis. This was followed by work of Baldwin and Lachlan [BL71] that showed that this result can be explained by finding—in a class of models of a countable theory that is categorical in an uncountable cardinal—admit a dimension invariant that determines the isomorphism class of the model. Shelah [She74] further extended Morley’s result showing that it also holds amongst uncountable theories, although the cutoff must be larger.
@@
Fact 3.1 (Morley [Mor65a], Baldwin-Lachlan [BL71], Shelah [She74]).
@@
In particular, we can take $\mu_{\lambda}=\lambda^{+}$. If $\lambda=\omega$, then there is a dimension associated with the models of $T$ that determines the isomorphism class (including for countable models). [^15] Unfortunately, Shelah’s extension to uncountable theories does not come with an associated dimension. This means that the motivating example from real vector spaces is not actually explainable by the abstract model theory. However, the linear dimension for vector spaces of countable field (such as $\mathbb{Q}$) or the transcendence dimension for algebraically closed fields of a fixed characteristic is discoverable via the Baldwin-Lachlan result).
@@
Conjecture 3.2 (Shelah’s Eventual Categoricity Conjecture).
@@
The standard version (with no ‘Eventual’) speicifies that $\mu_{\lambda}=\beth_{\left(2^{\lambda}\right)^{+}}$ (which is the Hanf number $H(\lambda)$), but even proving the more general eventual version stated above would be a huge advance.
@@
    Suppose that $2^{\kappa^{+n}}<2^{\kappa^{+(n+1)}}$ for all $n<\omega$. If $\mathbb{K}$ is categorical in $\kappa^{+n}$ for all $n<\omega$, then $\mathbb{K}$ is categorical in all $\kappa^{\prime}\geq\kappa$ ([She09c, Conclusion 12.43])
@@
    Suppose $\mathbb{K}$ has amalgamation, no maximal models and is $\kappa$-tame. Set $\mu_{\lambda}=\lambda^{+}+\kappa$. If $\mathbb{K}$ is categorical in $\kappa^{+}>\mu_{\lambda}$, then $\mathbb{K}$ is categorical in all $\kappa^{\prime}>\kappa^{+}$. ([GV06a, Theorem 5.2], Upward Categoricity from a Successor with Tameness)
@@
We should also comment on the description of Shelah’s Categoricity Conjecture as a ‘test question.’. This is used to indicate that it is not expected that a positive solution will lead to new results by citing the true statement. Instead, proving the conjecture is a test of the robustness of the structure theory of Abstract Elementary Classes, and it is the techniques used to solve it that will be used to prove other theorems. This has been seen in the partial results listed above, e.g., tameness in Fact 3.3.(3) or good $\lambda$-frames in Fact 3.3.(2).
@@ 3.1.2. Categoricity in Abstract Elementary Classes of Modules
We now situate categoricity in the study of Abstract Elementary Classes of modules. An important point is that categoricity is solely a property of the models in an Abstract Elementary Class $\mathbb{K}$ and is not impacted by the particular choice of strong substructure $\prec_{\mathbb{K}}$ (although some strong substructure must be given to witness that the collection of modules form an Abstract Elementary Classes). One theme throughout the results is that the threshold cardinals (which we normally denote $\mu_{\lambda}$) are much lower in the module examples than in general: compare the bound $|R|+\aleph_{0}$ in the clauses of Fact 3.6 below with the corresponding bound $\beth_{(|R|+\aleph_{0})^{+}}$ in Fact 3.3.(5) above.
@@
Given a module $M$ and a cardinal $\lambda$, we use $M^{(\lambda)}$ to denote the direct sum $\oplus_{\lambda}M$ of $\lambda$-many copies of $M$.
@@
In nice classes, there is a tight relation between categoricity and the number $\mu_{\mathbb{K}}$ of strongly indecomposable modules.
@@
There are many examples of classes of modules that turn out to be nice. In many cases, the categorical side of this dichotomoy (where $\mathbb{K}$ is categorical in some/all $\lambda>\mu_{\mathbb{K}}$) also comes with a characterization of the ring $R$. The various definitions of the ring properties are not crucial for our purposes, but are available at the citations.
@@
    This class is categorical in some power iff $R$ is the ring of matrices over some division ring. [MA23a, Lemma 2.12]
@@
    This class is categorical in some power iff $R$ is a full matric ring over a division ring (and thus reduces to the previous case). [MA23a, Theorem 3.4]
@@
    This class is categorical in some power iff $R$ is left Noetherian and there is a unique indecomposable injective module up to isomorphism. [MA23a, Theorem 3.11]
@@
    This class is categorical in some power iff $R$ is left Noetherian and there is a unique indecomposable injective module up to isomorphism. [MA23a, Theorem 3.19]
@@
    This class is categorical in some power iff $R$ is left perfect and $R/J(R)$ is the ring of matrices over some division ring iff $R$ is the ring of matrices over a local ring with left $T$-nilpotent maximal ideal. [MA23a, Theorem 3.26]
The extra information about $R$ in these cases comes from showing that relevant properties interact nicely with direct sums and any special module we care about is the direct sum of our strongly indecomposable generator. We give an example below. Recall that if every absolutely pure left $R$-module is injective, then $R$ is left Noetherian.
@@
Another direction [vT, Trl] for categoricity focuses on strong splitters (Definition 3.7.(3))as the special models in the context of deconstructible classes of modules.
@@
    The class $\mathbb{K}$ is deconstructible iff there is some $\kappa$ such that $\mathbb{K}$ is precisely the collection of all modules with a $\mathbb{K}_{*<\kappa}$-filtration, where $\mathbb{K}_{*<\kappa}$ are the $<\kappa$-presented modules in $\mathbb{K}$. In this case, we say $\mathbb{K}$ is generated by $\mathbb{K}_{*<\kappa}$.
@@
    for some $\kappa\geq|R|+\aleph_{0}$ such that $M$ is $\leq\kappa$-presented. [^17] Note the definition requires this condition for all direct sums, but [Trl, Lemma 2.2] shows that this single set is enough..
Note that $M$ being a strong splitter means that given any $N\supset M^{(\kappa)}$ such that $N/\left(M^{(\kappa)}\right)\cong M$, we must have a submodule $Q\subset N$ such that $M^{(\kappa)}\oplus Q=N$, that is, $M^{(\kappa)}$ splits in $N$.
Importantly, not all deconstructible classes of modules are Abstract Elementary Classes (implicitly using the submodule relation unless we state otherwise) and the following results cover all deconstructible classes. However, there are many examples of deconstructible classes that are Abstract Elementary Classes: any Roots of Ext (Example 2.8.(7)) that form an AEC are deconstructible [vT, Theorem 2.2].
In deconstructible classes, strong splitters are closely tied to categoricity. In fact, the existence of a single non-strong splitter massively restricts the categoricity spectrum. A helpful observation is that it suffices to check a particular model to see if there are any non-strong splitters.
@@
Fix a deconstructible class $\mathbb{K}$ that is generated by $\mathcal{C}$. Define
@@
Let $\mathbb{K}$ be a deconstructible class of $R$-modules. Suppose there is a non-strong splitter $M\in\mathbb{K}$ of size $\kappa\geq|R|+\aleph_{0}$. Then $\mathbb{K}$ is categorical in at most one cardinal $\lambda\geq\kappa$ and such a $\lambda$ is either equal to $\kappa$ or singular.
Of course, these restrictions on $\lambda$ beg the question if such a $\lambda$ could possibly exist. If $\mathbb{K}$ is $\mu^{+}$-deconstructible with $\mu\geq|R|+\aleph_{0}$, then such a $\lambda$ must be $\leq 2^{\mu}$ by [vT, Theorem 4.1]. This question is open [Trl, p. 381].
@@
To do this, build a continuous, increasing chain $\{M_{\alpha}:\alpha<\mu\}$ starting with $M_{0}=M^{(\lambda)}$ and using the following successor step: if $M_{\alpha}$ is isomorphic to some $\oplus_{\chi}M$ (necessarily with $\chi\geq\lambda$), then the definition of not strongly splitting means that there is an extension $M_{\alpha+1}\supset M_{\alpha}$ with $M_{\alpha+1}/M_{\alpha}\cong M$ and $M_{\alpha}$ does not split in $M_{\alpha+1}$. The first condition ensures $M_{\alpha+1}$ (and the eventual union) is in the deconstructible class, and the second condition will force us away from the union being a direct sum itself.
Set $M^{*}_{\mu}=\bigcup_{\alpha<\mu}M_{\alpha}$. The proof that $M^{*}_{\mu}$ is not isomorphic to $\oplus_{\mu}M$ is different in each case, but the essential argument is to a back and forth argument from a supposed bijection that contradicts our construction: it finds a stage $M_{\beta}\subset M_{\beta+1}$ such that $M_{\beta}$ is a direct summand of $M$ and, hence, $M_{\beta}$ splits in $M_{\beta+1}$. †
On the structure side, we may assume that all models are strong splitters. Then categoricity reduces to finding a power with all direct sums of generating modules to that power are isomorphic.
@@
Second, assume the right side and take some $N_{0}\in\mathcal{C}$ with $\|N_{0}\|\leq\nu<\lambda$. Fix $M\in\mathbb{K}_{\lambda}$ and we will show this is isomorphic to $N_{0}^{(\lambda)}$. Since every module in $\mathcal{C}$ is a strong splitter, we have that
@@
$M$$\cong$ $M_{0}\oplus M_{1}$
For $M_{0}$, we have $|\mathcal{C}_{0}|\leq|\mathcal{C}|<\lambda$, so $\|M_{0}\|<\lambda$ and $\|M_{1}\|=\lambda$. For $M_{1}$, we use our assumption and the fact $\kappa_{N}\geq\nu$ to conclude
@@
By our assumption, $M_{0}$ is isomorphic to a direct summand of $N_{0}^{(\lambda)}$ and $M\cong M_{0}\oplus N_{0}^{(\lambda)}$. So Eilenberg’s trick can be adapted (see the proof of [vT, Lemma 4.5]) to show